Fermi National Laboratory
 

Operation Summary for May 17, 2010 - May 19, 2010

 

Monday May 17

The day shift began with Operations monitoring the Tevatron (TeV) store 7816 (89.78E30) and the Recycler stash (404.7E10), with the Antiproton Source (Pbar) stacking (12.6E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, with beam to the Fermilab Test Beam Facility (FTBF), and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations disabled beam to FTBF at 8:20 AM for the installation of experiment T-1005. Operations terminated store 7816 at 9:40 AM. Operations brought the TeV to 150 GeV. Cryo system personnel performed wet engine repairs at TeV sector A4. Operations put the TeV into HEP shot setup at 11:45 AM and established store 7817 at 12:45 PM with an average initial luminosity (AIL) of 324E30. NuMI personnel raised the current limit for a Lambertson (LAM61) power supply to decrease the number of NuMI beam permit trips. Linac personnel swapped the phase controller in a quadrupole (Q5Bulk) power supply.

During the evening shift, Operations reported that the Pbar target pulsed magnet developed a noisy return LCW flow interlock readback, which was bypassed at the request of Pbar Target personnel. Operations received the Operational Readiness Clearance for the T-1005 experiment and delivered them beam for a brief period of time. Operations reported minimal downtime due to Linac RF problems.

Tuesday May 18

The midnight shift began with Operations monitoring store 7817 (83.54E30) and the stash (375.1E10), with Pbar stacking (9.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7817 at 12:30 AM, put the TeV into shot setup, and established store 7818 at 2 AM with an AIL of 371.9E30. The Booster RF west anode power supply tripped off at 5 AM, then after it was reset, BRF18 failed; this reduced the number of operational stations in Booster to 17. Operations established beam to T-1005 7 AM.

During the day shift, Booster RF personnel disconnected BRF18 from the anode supply. Linac personnel repaired a LCW leak on a Klystron RF station (KRF2). RF personnel investigated a MI low level RF VME crate that had developed an issue requiring hardware reboots.

During the evening shift, beam to NuMI was held off for about one hour due to problems with the hardware responsible for monitoring beam in the kicker gap; MI personnel power cycled the crate and restored the appropriate ACNET parameters. Operations disabled the T-1005 beam at 6:02 PM. Operations terminated store 7818 at 9:10 PM, put the TeV into shot setup, and established store 7819 at 10:20 PM with an AIL of 266.7E30.

Wednesday May 19

The midnight shift began with Operations monitoring store 7819 (229.49E30) and the stash (85.4E10), with Pbar stacking (8.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7819 aborted at 12:37 AM when a CIA crate at C2 lost power, which closed the sector beam valves due to a DC power supply failure. There was no quench. Operators called in Controls personnel to investigate the failure; they swapped power supplies. Operations turned on the TeV and put it into shot setup at 3:15 AM. Operations established store 7820 at 4:18 AM with an AIL of 187.7E30. Operations established beam to T-1005 at 4:30 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a two-hour access into CDF, D0, and the TeV after store 7820 is terminated later this morning.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/19/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory